OpenStudy (anonymous):

I have the function:\[f(x)=\sum_{n=1}^{\infty}\frac{ (-1)^{n+1} }{ 2n }x^n, x \in]-R,R[\] And I need to find the Maclaurin series for \[f´(x)\]I have found the convergence radius to 1. Now since R>0, then I can use term-by-tern differentiation right?

OpenStudy (anonymous):

Why do we need \(R>0\)?

OpenStudy (anonymous):

I wanted to use term-by-term differentiation. And that can only be done if we have a power series with radius of convergence \[R>0\]

OpenStudy (anonymous):

Right, you can use term-by-term differentiation.

OpenStudy (anonymous):

I am not 100% how to use that. Would that give me \[f'(x)=\sum_{n=2}^{\infty}\frac{ (n+1)(-1)^n }{ 2 }x^n\] @mukushla

OpenStudy (anonymous):

\((x^n)'=n x^{n-1}\) and you don't need to change the lower bound